### v3

# Spectral Flow Learning Theory: Finite-Sample Guarantees for Vector-Field Identification

Abstract
We study the identification of continuous-time vector fields from irregularly sampled trajectories. We introduce Spectral Flow Learning (SFL), which learns in a windowed flow space using a lag-linear label operator that aggregates lagged Koopman actions. We provide finite-sample high-probability (FS-HP) guarantees for the class of variable-step linear multistep methods (vLLM). The FS-HP rates are constructed using spectral regularization with qualification-controlled filters for flow predictors under standard source and filter assumptions. A multistep observability inequality links flow error to vector-field error and yields two-term bounds that combine a statistical rate with an explicit discretization bias from vLMM theory.
This preliminary preprint states the results and sketches proofs, with full proofs and extensions deferred to a journal version.
I Introduction
In this manuscript, we develop preliminary components of a Spectral Flow Learning (SFL) theory, a spectral-regularization view of learning dynamical laws from irregular trajectories. In SFL, learning is posed as an inverse problem [1] driven by the inclusion map $\mathsf{I}$ of a vector-valued Reproducing Kernel Hilbert Space (vvRKHS) into a population $L^{2}$ space; the covariance operator ${\mathsf{T}}={\mathsf{I}}^{*}{\mathsf{I}}$, a spectral filter ${\mathsf{g}}_{\lambda}$, its qualification $\nu$, and a Lipschitz exponent $\mu$ govern approximation/stability rates [2, 3].
We take the flow as the primary regression object and lift the finite-sample high-probability (FS-HP) excess risk bound from flow space to the vector field via an observability inequality induced by a Koopman-lag forcing operator $\mathsf{B}$. Our operator notation and bias-variance decomposition follow [4]. Deterministic modeling error from discretization is controlled by variable-step size linear multistep method (vLMM) theory, bounded coefficients, variable-step consistency, and ODE regularity condition [5, Ch. 3.5].
II Setup and Problem Formulation
This section makes the setup precise. We first fix notation and basic functional-analytic conventions. We then specify the dynamical system and its flows (autonomous and input-driven), introduce the design law that governs state visitation and the step law that governs time advances on irregular grids. Finally, we state the finite-sample, high-probability bounds to be proved later (Eq. (1)). These choices determine the population spaces and bounded operators used throughout.
II-A Notation
$\mathbb{R}$ denotes the real number line. Vectors in $\mathbb{R}^{n}$ are column vectors. For a set $A$, $\mathbf{1}_{A}$ denotes its indicator function. For a test function $\varphi$ and a measure $\mu$, the pairing is defined by $\langle\varphi,\mu\rangle\coloneqq\int\varphi d\mu.$ We write $a\lesssim b$ when there exists a universal constant $C>0$ with $a\leq Cb$, and we write $a\asymp b$ when $a\lesssim b$ and $b\lesssim a$. For a Hilbert space $\mathcal{H}$, $\langle u,v\rangle_{\mathcal{H}}$ denotes the inner product and $\|u\|_{\mathcal{H}}$ denotes the norm. On $\mathbb{R}^{n}$ the inner product is $\langle x,y\rangle=x^{\top}y$ and the Euclidean norm is $\|x\|_{2}=(x^{\top}x)^{1/2}$. The space $L^{2}(\mathcal{X},\mu)$ consists of square-integrable functions and its norm is $\|f\|_{L^{2}}=\sqrt{\int|f|^{2}d\mu}.$ The space $\mathcal{C}(\mathcal{X})$ denotes continuous real-valued functions on $\mathcal{X}$, and $\mathcal{C}^{k}(\mathcal{X})$ denotes the $k$-times continuously differentiable functions. The Banach space $\mathcal{L}(\mathcal{H}_{1},\mathcal{H}_{2})$ denotes bounded linear operators $\mathsf{T}:\mathcal{H}_{1}\to\mathcal{H}_{2}$, and the operator norm is $\|\mathsf{T}\|_{\mathcal{L}}=\sup_{\|x\|=1}\|\mathsf{T}x\|.$ The adjoint of $\mathsf{T}$ is written $\mathsf{T}^{*}$, and $\mathrm{Id}$ denotes the identity operator. An operator is self-adjoint when $\mathsf{T}=\mathsf{T}^{*}$, and it is positive semidefinite when $\mathsf{T}\succeq 0$. For a self-adjoint operator $\mathsf{T}$, $\lambda_{\min}(\mathsf{T})$ and $\lambda_{\max}(\mathsf{T})$ denote its minimal and maximal eigenvalues. The set $\mathcal{P}(\mathcal{X})$ denotes Borel probability measures on $\mathcal{X}$. For $\rho\in\mathcal{P}(\mathcal{X})$, $\operatorname{supp}(\rho)$ denotes the support and $\delta_{x}$ denotes the Dirac mass at $x$. We write $\mathbb{E}[\cdot]$ for expectation and $\mathbb{E}[\ \cdot\mid\cdot\ ]$ for conditional expectation. We write $a_{n}=\mathcal{O}(b_{n})$ when $\sup_{n}|a_{n}/b_{n}|<\infty$.
II-B Dynamical System and Flows
In SFL, the flow is the primary regression object: we view the problem of learning vector-fields from trajectory data as identifying the one–step map $(x,h)\mapsto\Phi^{h}(x)$ in an $L^{2}$ space and then lift the flow estimator to the vector field. This choice makes the feature-label pairs well-posed under irregular sampling and enables us to apply spectral regularization methods in learning theory to system identification problems. With this motivation, we now define the dynamical model and its flows. Let $\mathcal{X}\subset\mathbb{R}^{n}$ be compact and $\mathcal{B}=\mathcal{B}(\mathcal{X})$ its Borel $\sigma$-algebra. Consider the control‑affine dynamics:
$\dot{x}(t)=f(x(t))+g(x(t))u(t),\qquad x(0)=x_{0},$
with $f:\mathcal{X}\to\mathbb{R}^{n}$, $g:\mathcal{X}\to\mathbb{R}^{n\times m}$ locally Lipschitz, and $u:\mathbb{R}_{\geq 0}\to\mathbb{R}^{m}$ measurable, locally bounded. Assume forward invariance: $x(t)\in\mathcal{X}$ for $t\geq 0$.
Furthermore, let autonomous flow with zero input $\phi^{t}:\mathcal{X}\to\mathcal{X}$ solve $\tfrac{d}{dt}\phi^{t}(x)=f(\phi^{t}(x))$ with $\phi^{0}=\mathrm{id}$ and semigroup property $\phi^{t+s}=\phi^{t}\circ\phi^{s}$. Define input-driven flow $\Phi_{u}^{t,s}:\mathcal{X}\to\mathcal{X}$ that maps $x(s)\mapsto x(t)$ and satisfies:
$\tfrac{d}{dt}\Phi_{u}^{t,s}(x)$ $=f(\Phi_{u}^{t,s}(x))+g(\Phi_{u}^{t,s}(x))u(t),$
$\Phi_{u}^{s,s}$ $=\mathrm{id},\quad\Phi_{u}^{t,s}=\Phi_{\tau^{s}u}^{t-s,0}.$
where $\tau^{s}u$ is the time-shift operator of the input signal $u$ by $s$, such that $(\tau^{s}u)(t)\coloneqq u(t+s),\forall t\geq 0$. We write $\Phi_{u}^{t}\coloneqq\Phi_{u}^{t,0}$. Special case: $u\equiv 0\Rightarrow\Phi_{0}^{t}=\phi^{t}$. We denote flows generated by an autonomous vector field $f$ and control-affine dynamics $(f,g)$ as $\phi^{h}(\ \cdot\ ;f)$ and $\Phi^{h}_{u}(\ \cdot\ ;f,g)$ respectively. Furthermore, we use $\Phi^{h}(\cdot)$ when the discussion is agnostic with respect to the existence of a system input $u$, and $\Phi^{h}_{\rho}(\cdot)$ when the input is determined by a design law $\rho$.
II-C Design Law (state visitation distribution)
The design law $\rho_{\mathcal{X}}^{\mathrm{design}}\in\mathcal{P}(\mathcal{X})$, specifies where the experiment spends time in $\mathcal{X}$. $\rho_{\mathcal{X}}^{\mathrm{design}}$ is a Borel probability measure on $(\mathcal{X},\mathcal{B})$. There are multiple ways to construct a design law; we are interested in the continuous-time time-average law. Along a single path $x(t)=\Phi_{u}^{t}(x_{0})$:
$\lim_{T\to\infty}\bigg\langle\varphi,\frac{1}{T}\int_{0}^{T}\delta_{x(t)}dt\bigg\rangle=\langle\varphi,\rho_{\mathcal{X}}\rangle,\quad\forall\ \varphi\in\mathcal{C}(\mathcal{X}).$
where $\delta_{x}$ is the Dirac measure: $\delta_{x}(A)=\mathbf{1}_{A}(x)$, $A\in\mathcal{B}$. The design law $\rho_{\mathcal{X}}^{\mathrm{design}}$ determines coverage of states; identifiability over the identificaiton domain $\mathcal{X}_{\rm ID}$ requires $\mathcal{X}_{\rm ID}\subseteq\operatorname{supp}(\rho_{\mathcal{X}}^{\mathrm{design}})$.
Remark II.1.
The design law induces the geometry in which we measure errors and therefore governs identifiability. In particular, we can view identifiability as an input–design problem: choose an input policy $u(\cdot)$ so that the induced $\rho_{\mathcal{X}}^{\mathrm{design}}(u)$ places mass on the domain of interest. Moreover, the population risks $\mathcal{E}(\Phi_{\rho})$ and $\mathcal{E}(f_{\rho})$ used later in the problem formulation are squared $L^{2}$-errors with respect to the design weights $\rho^{\rm design}_{\mathcal{X}}\times\rho_{\mathbf{h}}$ and $\rho^{\rm design}_{\mathcal{X}}$. The requirement of $\mathcal{X}_{\rm ID}\subseteq\operatorname{supp}(\rho_{\mathcal{X}}^{\mathrm{design}})$ ensures that discrepancies on the identification domain are penalized and thus identifiable.
II-D Step Law (time‑advance distribution)
Let $\boldsymbol{h}\coloneqq(0,h_{\max}]$, the step law $\rho_{\boldsymbol{h}}\in\mathcal{P}\big((0,h_{\max}]\big)$ specifies how far we advance in time when forming one‑step labels $(x_{k},h_{k})\mapsto x_{k+1}$. It is a probability measure on the interval $(0,h_{\max}]$. In this work, we are interested in a deterministic schedule $\{h_{k}\}$. Specifically, a step law $\rho_{\boldsymbol{h}}$ can be constructed through index‑average:
$\lim_{N\to\infty}\bigg\langle\varphi,\frac{1}{N}\sum_{k=0}^{N-1}\delta_{h_{k}}\bigg\rangle=\langle\varphi,\rho_{\boldsymbol{h}}\rangle,\quad\forall\ \varphi\in\mathcal{C}(\mathcal{X}).$
We are ready to formulate the problem of identifying flow and vector field FS-HP bounds in the SFL framework.
II-E Problem Formulation (FS-HP bound)
Fix the design law $\rho_{\mathcal{X}}^{\mathrm{design}}\in\mathcal{P}(\mathcal{X})$ and the step law $\rho_{\boldsymbol{h}}\in\mathcal{P}((0,h_{\max}])$. Let $\ell$ denote the number of effective samples and let $H\sim\rho_{\boldsymbol{h}}$ denote the anchor step, where $\mathbb{E}[H^{q}]$ reduces to $h^{q}$ for a deterministic schedule. We define the flow feature-label pairs $\{((x_{i},h_{i}),y_{i})\}_{i=1}^{\ell}$. Let $\rho((x,h),y)=\rho(y\mid x,h)\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}(x,h)$ with $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}\coloneqq\rho_{\mathcal{X}}^{\mathrm{design}}\times\rho_{\boldsymbol{h}}$. We denote the population flow space as $\mathcal{W}$ and likewise the population vector-field space as $\mathcal{V}$. The target flow $\Phi_{\rho}\in\mathcal{W}$ is:
$\Phi^{h}_{\rho}(x)\coloneqq\mathbb{E}[Y\mid X=x,H=h]=\int_{\mathcal{X}}yd\rho(y\mid x,h).$
Similarly, the target vector-field $f_{\rho}\in\mathcal{V}$ is defined as:
$f_{\rho}(x)\coloneqq\lim_{r\downarrow 0}\;\mathbb{E}\!\left[\frac{Y-X}{H}\bigg|X=x,0<H\leq r\right],$
Under the flow source condition $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$, [4, Eq. 9], and a spectral filter ${\mathsf{g}}_{\lambda}$ with qualification $\nu\geq r$ and Lipschitz exponent $\mu$ (set $\beta\coloneqq\max\{1,2\mu\}$), [4, Def. 1], our goal is to establish finite-sample, high-probability (FS–HP) bounds of the form:
$\underbrace{\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}}_{\text{flow excess risk}}$ $\leq\log\!\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}},$ (1)
$\underbrace{\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}}_{\text{vector-field excess risk}}$ $\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}}\log\!\frac{4}{\eta}\ell^{-\frac{2r}{2r+\beta}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}}\mathbb{E}[H^{2p+2}],$ (2)
which holds with probability at least $1-\eta$ over the model $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}$. Fig. 1 shows the road-map to proof the FS-HP bounds for the flow excess risk in Theorem 1 and the FS-HP bounds for the vector field excess risk in Theorem 2.
vLMM Assumptions Spectral Regularization Asm. III.1 bounded step ratios/coeffs ($\zeta,\alpha,\beta$) Filter qual./Lipschitz (Def. [4, Def. 1]) Asm. III.2 order-$p$ consistency Source cond. on vector field $\Phi_{\rho}\in\Omega^{(\mathcal{V})}_{r,R}$ Asm. III.3 ODE regularity Source cond. on flow $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ Lemma 2 Uniform LTE Lemma 1 Observability via $\mathsf{B}$ Thm. 1 Flow FS–HP bound Thm. 2 Vector-field FS–HP bound
Figure 1: Dependency graph linking assumptions, lemmas, theorems, corollaries, and propositions used in the SFL analysis.
III Variable Step Linear Multistep Method (vLMM) Theory
This section records some vLMM facts we use to turn flow-space error into vector-field error within SFL. The role of vLMM in the SFL framework is twofold: (i) it supplies the windowed surrogate that links vector fields to flows through the multistep forcing/aggregation operators (enabling observability), and (ii) it controls the deterministic discretization bias via order-$p$ truncation guarantees. Since, for variable-step methods, consistency depends only on step-ratio and not on absolute time shifts, a step-window is often parameterized by the step-ratio vector $\zeta$ and an anchor step $h$ in the vLMM framework.
III-A General Form
Let there be a context window of length $M$ with step-window $h_{1},\dots,h_{M-1}]\in\boldsymbol{h}$. Denote the future step as the window anchor, $h\coloneqq h_{M}$. The bounded step-ratio vector is defined as:
$\zeta\coloneqq[\zeta_{1},\dots,\zeta_{M-1}]\in\boldsymbol{\zeta}\subset(0,\bar{\zeta}]^{M-1},$
where $\zeta_{k}\coloneqq h_{k}/h_{k-1}$ and $\zeta_{M-1}=h/h_{M-1}$.
Definition 1 (vLMM General Form).
A vLMM takes the following general form [5, Eq. 5.15]:
$x_{k+M}+\sum_{j=1}^{M-1}\alpha_{jk}x_{k+j}=h_{k+M-1}\sum_{j=1}^{M}\beta_{jk}f(x_{k+j}),$ (3)
with coefficients: $\alpha_{jk}\coloneqq\alpha_{j}(\zeta_{k+1},\dots,\zeta_{k+M-1}),\beta_{jk}\coloneqq\beta_{j}(\zeta_{k+1},\dots,\zeta_{k+M-1}),$ functions of the step ratios $\zeta$.
Before stating the deterministic bias assumptions, we instantiate the unified form (3) with the standard variable-step families: Adams-Bashforth (AB), Adams-Moulton (AM), and backward-differentiation formulas (BDF) [5]. Each of the above method arises from a suitable choice of the ratio-dependent coefficients, $(\alpha_{jk}(\zeta),\beta_{jk}(\zeta))$, and admits a windowed-flow representation aligned with our SFL label operator, bringing explicit, implicit, and stiff-stable linear discretizations under one single framework. Furthermore, notice that the popular Savitzky-Golay (SG) derivative estimators arise from local polynomial least-squares fitting on a sliding window [6, 7]. Therefore, for uniform steps and one-sided windows, SG estimators reduce to the classical finite-difference derivative weights underlying multistep formulas. Within SFL framework, SG estimators can be seen as an alternative windowed label operator providing derivative labels complementary to vLMM flow labels.
III-B Assumptions for Deterministic Bias
We now list the assumptions needed to obtain deterministic bounds on the vLMM-induced flow approximation error.
Definition 2 (Order-$p$ consistency).
A vLMM (3) is order-$p$  [5, Def. 5.2] if:
$q(x_{k_{M}})+\sum_{j=1}^{M-1}\alpha_{jk}q(x_{k+j})=h_{k+M-1}\sum_{j=0}^{M}\beta_{jk}q^{\prime}(x_{k+j})$
holds for all polynomials $q(x)$ of degree $\leq p$ and grids $(h_{j})$.
Notice that the window size $M$ is the number of back steps a $M$-step vLMM uses, whereas the order $p$ is the accuracy of the method. They are related but not identical: in classical families, AB has $p=M$, AM has $p=M+1$, and BDF has $p=M$ (with zero-stability only up to $M\leq 6$). Moving onward, we state the key assumptions needed from the vLMM theory to establish an FS-HP bound in the SFL framework.
Assumption III.1 (Bounded step ratios and coefficients).
There exist constants $0<\underline{\zeta}\leq\bar{\zeta}<\infty$ such that:
$\underline{\zeta}\ \leq\ \zeta_{k}\ \leq\ \bar{\zeta}\qquad\forall k.$
Moreover, the variable–step coefficients $\{\alpha_{jk},\beta_{jk}\}$ are uniformly bounded.
Notice that bounded coefficients $\alpha_{jk},\beta_{jk}$ follow from bounded step ratios for AB/AM methods, and BDF [5, Lem. 5.3].
Assumption III.2.
The vLMM (3) is order-$p$ consistent in the sense of Definition 2.
Assumption III.3 (ODE regularity).
The right–hand side $f$ is sufficiently smooth, i.e., $f\in\mathcal{C}^{p+1}$ and locally Lipschitz on the forward–invariant compact set $\mathcal{X}$.
IV Spectral Regularization Learning Theory
This section assembles the spectral-regularization ingredients used by SFL. We adopt the operator-theoretic viewpoint of [4, 2, 3]: population and sampling Hilbert spaces are linked by inclusion maps whose covariance operators are spectrally filtered to stabilize inversion. Concretely, we fix the vector-field space $\mathcal{V}$, the population flow space $\mathcal{W}^{\circ}$, and the windowed flow space $\mathcal{W}$ on which learning acts; define the inclusions $\mathsf{I}_{\mathcal{V}},\mathsf{I}_{\mathcal{W}}$ with covariance operators $\mathsf{T}_{\mathcal{V}},\mathsf{T}_{\mathcal{W}}$ (and integral operators $\mathsf{L}_{\mathcal{V}},\mathsf{L}_{\mathcal{W}}$); and introduce a spectral filter ${\mathsf{g}}_{\lambda}$ with qualification $\nu$ and Lipschitz exponent $\mu$. In SFL, the estimator is obtained by filtering $\mathsf{T}_{\mathcal{W}}$ (componentwise in the vvRKHS), yielding a flow predictor in $\mathcal{H}_{\mathcal{W}}$ that we subsequently lift to the vector field using the multistep observability inequality and vLMM bias control on irregular grids.
IV-A Population and Sampling Hilbert Spaces
We first encounter the notion of population flow $\mathcal{W}$ and vector-field space $\mathcal{V}$ in the problem formulation. In this section, we refine the two notions by formally introducing the three population Hilbert spaces $\mathcal{V},\mathcal{W},\mathcal{W}^{\circ}$ and a finite-dimensional sampling space $\mathcal{Y}$. The population vector-field space $\mathcal{V}$ is defined as:
$\mathcal{V}\coloneqq L^{2}(\mathcal{X},\rho_{\mathcal{X}}^{\rm design};\mathbb{R}^{n}),\quad\mathcal{H}_{\mathcal{V}}\subset\mathcal{V},$
where $\mathcal{H}_{\mathcal{V}}$ is a vvRKHS equipped with an inner product, $\langle\cdot,\cdot\rangle_{\mathcal{H}_{\mathcal{V}}}$, induced by the measurable, bounded PHS kernel $k_{\rm phs}$ and densely embedded in $\mathcal{V}$ with the inclusion map $\mathsf{I}_{\mathcal{V}}:\mathcal{H}_{\mathcal{V}}\to\mathcal{V}$. Recall that $\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}\coloneqq\rho_{\mathcal{X}}^{\mathrm{design}}\times\rho_{\boldsymbol{h}}$, the single-step flow space is:
$\mathcal{W}^{\circ}\coloneqq L^{2}\big(\mathcal{X}\times\boldsymbol{h},\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design};\mathbb{R}^{n}\big),\quad\mathcal{H}_{\mathcal{W}^{\circ}}\subset\mathcal{W}^{\circ},$
with the embedded vvRKHS $\mathcal{H}_{\mathcal{W}^{\circ}}$ and the inclusion map defined as $\mathsf{I}_{\mathcal{W}^{\circ}}:\mathcal{H}_{\mathcal{W}^{\circ}}\to\mathcal{W}^{\circ}$. Notice that $\mathcal{W}^{\circ}$ is the space in which the one-step exact flow $\Phi^{h}(x;f_{\rho})$ lives. For the windowed flow space, recall that the window anchor is $h_{\rm anc}\in h$ and the bounded step-ratio vector is:
$\zeta\coloneqq[\zeta_{1},\dots,\zeta_{M-1}]\in\boldsymbol{\zeta}\subset(0,\bar{\zeta}]^{M-1},$
where $\zeta_{k}\coloneqq h_{k}/h_{k-1}$ and $\zeta_{M-1}=h_{\rm anc}/h_{M-1}$. Denote the backshift times by $\tau_{j}(\zeta)\geq 0$ measured from $h_{\rm anc}$ and a composite measure $\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}}\coloneqq\rho_{\mathcal{X}}^{\rm design}\times\rho_{\boldsymbol{h}}\times\rho_{\boldsymbol{\zeta}}$. Define the windowed flow space:
$\mathcal{W}\coloneqq L^{2}\big(\mathcal{X}\times\boldsymbol{h}\times\boldsymbol{\zeta},\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}};\ \mathbb{R}^{n}\big),\quad\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}.$
This is the space in which spectral regularization acts directly.
Notice that $\mathcal{W}^{\circ}$ can be embedded in $\mathcal{W}$ using the isometric lifting operator $\mathsf{J}:\mathcal{W}^{\circ}\to{\mathcal{W}}$. For every one-step flow $\Phi\in\mathcal{W}^{\circ}$, there is a $\mathsf{J}\Phi\in\mathcal{W}$:
$(\mathsf{J}\Phi)(x,h,\zeta)\coloneqq\Phi^{h}(x),$
that is, $\mathsf{J}\Phi$ is independent of $\zeta$ ($\zeta$-constant). An exact flow $\Phi^{h}(x;f)\in\mathcal{W}^{\circ}$ induced from $f$ is included in $\mathcal{W}$ via:
$\Phi^{h}(x;f)=(\mathsf{J}\Phi)(x,h,\zeta;f)\eqqcolon\Psi(x,h,\zeta;f)\in\mathcal{W}.$
Finally, the sampling space for finite data is $\mathcal{Y}\subset\mathbb{R}^{n\ell}$, equipped with the empirical norm $\|y\|_{\mathcal{Y}}^{2}=\frac{1}{\ell}\sum_{i=1}^{\ell}\|y_{i}\|^{2}$, with $\ell$ denoting the number of effective data points.
IV-B Inclusion Maps and Other Important Operators
Denote the inclusion $\mathsf{I}_{\mathcal{V}}:\mathcal{H}_{\mathcal{V}}\to\mathcal{V}$. Define the covariance operator $\mathsf{T}_{\mathcal{V}}$ and integral operator $\mathsf{L}_{\mathcal{V}}$:
$\mathsf{T}_{\mathcal{V}}\coloneqq\mathsf{I}_{\mathcal{V}}^{*}\mathsf{I}_{\mathcal{V}}:\mathcal{H}_{\mathcal{V}}\to\mathcal{H}_{\mathcal{V}},\qquad\mathsf{L}_{\mathcal{V}}\coloneqq\mathsf{I}_{\mathcal{V}}\mathsf{I}_{\mathcal{V}}^{*}:\mathcal{V}\to\mathcal{V}.$ (4)
For learning in flow space we fix a single hypothesis vvRKHS $\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}$ with bounded evaluation and inclusion $\mathsf{I}_{\mathcal{W}}:\mathcal{H}_{\mathcal{W}}\to\mathcal{W}$. Define:
$\mathsf{T}_{\mathcal{W}}\coloneqq\mathsf{I}_{\mathcal{W}}^{*}\mathsf{I}_{\mathcal{W}}:\mathcal{H}_{\mathcal{W}}\to\mathcal{H}_{\mathcal{W}},\qquad\mathsf{L}_{\mathcal{W}}\coloneqq\mathsf{I}_{\mathcal{W}}\mathsf{I}_{\mathcal{W}}^{*}:\mathcal{W}\to\mathcal{W}.$ (5)
We use spectral‑filter estimators $\Phi^{\lambda_{\ell}}_{z,y}\in\mathcal{H}_{\mathcal{W}}$ based on a filter ${\mathsf{g}}_{\lambda}$ applied to $\mathsf{T}_{\mathcal{W}}$ (vector‑valued extension is componentwise).
IV-C Filter Qualification and Lipschitz Condition
Another component we need for constructing a FS-HP bound is the notion of a spectral filter, its qualification, and the Lipschitz condition.
Definition 3.
[4, Def. 1] Let ${\mathsf{g}}_{\lambda}:[0,\kappa^{2}]\to\mathbb{R}$ be a spectral filter. Its qualification $\nu>0$ is the largest order such that:
$\sup_{0<\sigma\leq\kappa^{2}}\big|1-\sigma{\mathsf{g}}_{\lambda}(\sigma)\big|\sigma^{\nu}\ \leq\ \gamma_{\nu}\lambda^{\nu},\qquad\forall 0<\nu\leq\bar{\nu}.$
We also use the Lipschitz condition $|{\mathsf{g}}_{\lambda}(\sigma)-{\mathsf{g}}_{\lambda}(\sigma^{\prime})|\leq L\lambda^{-\mu}|\sigma-\sigma^{\prime}|$ for some $\mu>0$, and set $\beta=\max\{1,2\mu\}$.
A spectral filter is an approximate inverse applied to the spectrum of a positive self-adjoint operator (e.g., the covariance/integral operator) via functional calculus: if a covariance operator $\mathsf{T}=\sum_{i}\sigma_{i}\langle\cdot,e_{i}\rangle e_{i}$, then ${\mathsf{g}}_{\lambda}(\mathsf{T})=\sum_{i}{\mathsf{g}}_{\lambda}(\sigma_{i})\langle\cdot,e_{i}\rangle e_{i}$. It attenuates directions with small eigenvalues $\sigma_{i}$ to stabilize the inversion, while the residual $1-\sigma{\mathsf{g}}_{\lambda}(\sigma)$ quantifies the bias introduced by regularization. The qualification $\nu$ is the largest order for which $r_{\lambda}(\sigma)\leq\lambda^{\nu}\sigma^{-\nu}$ uniformly. Typical values of $\nu$ are: Tikhonov with $\nu=1$; order-$t$ iterated Tikhonov with $\nu=t$; Landweber with $\nu\to\infty$ [4, Sec. 3.1].
Consequently, under a source condition of order $r\leq\nu$ one obtains approximation error of order $\lambda^{2r}$ (in the prediction norm controlled by $\mathsf{T}$). The Lipschitz condition ensures stability of ${\mathsf{g}}_{\lambda}(\mathsf{T})$ to empirical perturbations $\mathsf{T}\mapsto\mathsf{S}_{\ell}\circ\mathsf{T}$, which is crucial for finite-sample FS-HP bounds. Gaussian process regression can be seen as Tikhonov regularization with $\lambda=\sigma_{y}^{2}/\ell$, with $\sigma_{y}^{2}$ the variance of observation noise.
IV-D Source Condition
In the previous section, we fix the hypothesis spaces $(\mathcal{H}_{\mathcal{V}},\mathcal{H}_{\mathcal{W}})$ and the spectral filter ${\mathsf{g}}_{\lambda}$, with qualification $\nu$ and Lipschitz exponent $\mu$. Moving onward, we quantify target-kernel alignment by a Hölder-type source condition [1] with exponent $r>0$; larger $r$ yields faster rates up to the filter’s qualification $\nu$. Conceptually, the source condition measures how well the targets $\Phi_{\rho}$ and $f_{\rho}$ align with the vvRKHS prior; formally, it requires them to lie in the range of a positive power of the integral operator $\mathsf{L}=\mathsf{I}\mathsf{I}^{*}$, and its vector-field analogue.
SFL learns dynamics in a windowed flow space and lifts both estimates and their guarantees to the vector field. Accordingly, we impose source conditions at both levels: a flow condition that directly drives the flow FS-HP bound, and a vector-field condition that encodes differential physics, e.g., Hamiltonian passivity, incompressibility $\nabla\cdot f=0$, and symmetry/equivariance. In practice, these constraints can be enforced via a physics-informed vvRKHS kernel [8], by projecting the lifted estimator onto the constraint set [9], or by adding a physics-informed penalty to the loss [10].
For $r>0$ and $R_{\mathcal{V}}>0$, assume $f_{\rho}$ is the target vector field with the following structure defined through (4):
$f_{\rho}\in\Omega^{(\mathcal{V})}_{r,R_{\mathcal{V}}}\ \coloneqq\ \big\{f\in\mathcal{V}:f=\mathsf{L}_{\mathcal{V}}^{r}v,\ \|v\|_{\mathcal{V}}\leq R_{\mathcal{V}}\big\}.$
Likewise, for $R_{\mathcal{W}}>0$,
$\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R_{\mathcal{W}}}\ \coloneqq\ \big\{\Phi\in\mathcal{W}:\ \Phi=\mathsf{L}_{\mathcal{W}}^{r}w,\ \|w\|_{\mathcal{W}}\leq R_{\mathcal{W}}\big\},$
where $\mathsf{L}_{\mathcal{W}}$ is define in (5). This flow-level source is the one used in Theorem 1 to obtain the FS-HP rate $\ell^{-\frac{2r}{2r+\beta}}$ (with $\beta=\max\{1,2\mu\}$).
Remark IV.1 (Compatibility/transferring source between levels).
If one wishes to derive the flow source from the vector-field source, it suffices to choose $\mathcal{H}_{\mathcal{W}}$ compatible with the exact flow labeling map $\mathsf{E}_{\rm exact}:\mathcal{H}_{\mathcal{V}}\to\mathcal{W}^{\circ}$ and the lifting $\mathsf{J}:\mathcal{W}^{\circ}\to\mathcal{W}$. Here $\mathsf{E}_{\rm exact}$ is the exact one-step flow map that sends a vector field to its true flow: $(\mathsf{E}_{\rm exact}f_{\rho})(x,h)\coloneqq\Phi^{h}(x;f_{\rho})$, and composing with $\mathsf{J}$ yields the $\zeta$-constant embedding $(\mathsf{J}\mathsf{E}_{\rm exact}f_{\rho})(x,h,\zeta)=\Phi^{h}(x;f_{\rho})$. Under this design, there is a constant $C_{\rm push}$ (depending only on operator norms) such that:
$f_{\rho}\in\Omega^{(\mathcal{V})}_{r,R_{\mathcal{V}}}\quad\Longrightarrow\quad\Phi_{\rho}=\mathsf{J}\mathsf{E}_{\rm exact}f_{\rho}\ \in\ \Omega^{(\mathcal{W})}_{r,R_{\mathcal{W}}}$
with $R_{\mathcal{W}}\leq C_{\rm push}R_{\mathcal{V}}$. In our general theory, we assume the flow source directly (to remain agnostic about the precise hypothesis design in $\mathcal{W}$), but the construction above shows how to enforce it by design.
V Spectral Flow Learning Theory
This section states the core constructions and guarantees of SFL. We formalize how variable-step multistep structure induces well-posed labels via the Koopman-lag forcing and state-aggregation operators, define the label map and vLMM residual, and provide an observability inequality (with constant $c_{\rm obs}$) together with a uniform local-truncation bound. These ingredients combine with spectral regularization in the windowed flow space to yield FS-HP-style, excess-risk bounds in the flow space and, via observability, in the vector-field space.
V-A Label Map and vLMM Residual
A key aspect of a theory on learning a vector field from trajectory data is the construction of well-defined input-label pairs. In our framework, we leverage the Koopman-lag forcing operator and state-aggregation operator to define the population level loss through the label map and vLMM residual.
Consider a vLMM as defined in Definition 1. Fix the window and anchor it at $t_{k+M}$. Define the backward lags:
$\tau_{0}:=0,\qquad\tau_{i}:=\sum_{q=1}^{i}h_{k+M-q}\quad(i=1,\dots,M),$
so that each state in the window satisfies:
$x_{k+M-i}=\Phi^{-\tau_{i}}\!\big(x_{k+M}\big),\qquad i=0,\dots,M.$
Equivalently, for $j=1,\dots,M$,
$x_{k+j}=\Phi^{-\tau_{M-j}}\big(x_{k+M}\big).$
Using the Koopman action $(U^{t}f)(x)=f(\Phi^{t}(x))$, the right-hand side of (3) can be rewritten as:
$h_{k+M-1}\sum_{j=1}^{M}\beta_{jk}(U^{-\tau_{M-j}}f)\big(x_{k+M}\big).$
Renaming the indices as $\beta_{j}(\zeta)\coloneqq\beta_{M-j,k}$ and setting $x\coloneqq x_{k+M}$, we define the Koopman-lag forcing operator:
$(\mathsf{B}f)(x;h,\zeta)\coloneqq h\sum_{j=0}^{M}\beta_{j}(\zeta)(U^{-\tau_{j}}f)(x)\quad:\ \mathcal{V}\to\mathcal{W},$ (6)
where $\tau_{j}(\zeta)$ are the cumulative lags determined by $\zeta$ and the anchor step $h\coloneqq h_{k+M-1}$. Similarly, the state-aggregation operator is defined as:
$(\mathsf{A}\Phi)(x;h,\zeta)\coloneqq\sum_{j=1}^{M}\alpha_{j}(\zeta)\Phi^{-\tau_{j}}(x)\quad:\ \mathcal{W}^{\circ}\to\mathcal{W}.$
Now, we are ready to define the population level label map and vLMM residual.
Definition 4 (Label Map and vLMM residual).
The population label map and vLMM residual are:
$\mathsf{Y}_{\alpha}(\Phi)\coloneqq\mathsf{J}\Phi+\mathsf{A}\Phi\in\mathcal{W},\ \Delta_{\mathrm{MS}}(f)\coloneqq\mathsf{Y}_{\alpha}(\Phi)-\mathsf{B}f\ \in\mathcal{W},$ (7)
respectively.
V-B Observability Inequality and Local Truncation Error (LTE)
Some useful facts we can leverage from the above construction include the following.
Lemma 1 (Observability via $\mathsf{B}$).
Let:
$c_{\rm obs}\coloneqq\inf_{\|v\|_{\mathcal{V}}=1}\|\mathsf{B}v\|_{\mathcal{W}}=\sqrt{\lambda_{\min}(\mathsf{B}^{*}\mathsf{B})},$
then:
$\|\mathsf{B}(f-f_{\rho})\|_{\mathcal{W}}^{2}\geq c_{\rm obs}^{2}\|f-f_{\rho}\|_{\mathcal{V}}^{2},\qquad\forall f\in\mathcal{V}.$
Proof.
This is immediate from $\|Av\|^{2}\geq\lambda_{\min}(A^{*}A)\|v\|^{2}$ for any bounded operator $A$. ∎
Lemma 2 (Uniform LTE for vLMM residual).
Under Assumptions III.1, III.2, and III.3, there exists a constant $C_{\mathrm{LTE},f}>0$, independent of $t$, $h$, and $\zeta$, such that the single-window residual is:
$\|\Delta_{\mathrm{MS}}(f)(x(t),h,\zeta)\|_{2}\leq C_{\mathrm{LTE},f}h^{p+1},$
for all $t\geq 0$, $h\in\boldsymbol{h}$, and $\zeta\in\boldsymbol{\zeta}$. Consequently, in the flow space $L^{2}$ norm,
$\|\Delta_{\mathrm{MS}}(f)\|_{\mathcal{W}}\leq\ C_{\mathrm{LTE},f}\sqrt{\mathbb{E}[H^{2p+2}]}.$ (8)
Proof.
Let $x(\cdot)$ solve $\dot{x}=x(t)=f(x(t))+g(x(t))u(t)$ on a forward‑invariant compact set $\mathcal{X}$ so that $\sup_{t}\|x^{(p+1)}(t)\|<\infty$. Following the notation in [5], we write the residual componentwise as the variable‑step difference operator:
$\mathcal{L}_{\zeta,h}[x]=x(t{+}h)$
$\quad+\sum_{j=1}^{M-1}\alpha_{j}(\zeta)x(t{-}\tau_{j}(\zeta))-h\sum_{j=0}^{M}\beta_{j}(\zeta)x^{\prime}(t{-}\tau_{j}(\zeta)).$
By Assumption III.2, which provides order-$p$ consistency, Assumption III.1 that guarantees bounded variable‑step coefficients, and Assumption III.3 for smooth ODE, [5, Eq. 5.17] yields:
$\mathcal{L}_{\zeta,h}[x]=\mathcal{O}(h^{p+1})\qquad\text{for smooth }x(\cdot),$
uniformly over admissible $\zeta$. A trivial extension of [5, Lem. 2.2] (constant‑step local error equivalence) to variable steps then implies that the local error at the anchor point is $\mathcal{O}(h^{p+1})$. Identifying $\Delta_{\mathrm{MS}}(f)=\mathcal{L}_{\zeta,h}[x]$ gives the pointwise bound with a constant $C_{\mathrm{LTE},f}$ independent of $(t,h,\zeta)$. This proves the first claim.
For the $\mathcal{W}$ norm, square and integrate over $\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}}$ to get:
$\|\Delta_{\mathrm{MS}}(f)\|_{\mathcal{W}}^{2}$ $=\int\|\Delta_{\mathrm{MS}}(f)(x,h,\zeta)\|^{2}_{2}\ d\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}}$
$\leq\ C_{\mathrm{LTE},f}^{2}\int h^{2p+2}d\rho_{\boldsymbol{h}}=C_{\mathrm{LTE},f}^{2}\mathbb{E}[H^{2p+2}].$
Taking the square root on both sides yields (8). ∎
Lemmas 1 and 2 provide stepping stones needed to lift flow-space guarantees to the vector field: a $\mathsf{B}$-observability inequality that controls $\|f-f_{\rho}\|_{\mathcal{V}}$ by $\|\mathsf{B}(f-f_{\rho})\|_{\mathcal{W}}$, and a uniform bound of the vLMM modeling residual via the moment $\mathbb{E}[H^{2p+2}]$. Next, we formalize the flow regression problem, introduce the excess risk $\mathcal{E}(\Phi)$ and the target regression function $\Phi_{\rho}$, and state the flow FS–HP bound obtained by spectral regularization. Combining that bound with the bounded label operator $y_{\alpha}=\mathsf{J}+\mathsf{A}$ and the decomposition $\mathsf{B}(\widehat{f}-f_{\rho})=\Delta_{\rm MS}(\widehat{f})+{\mathsf{Y}}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})+\Delta_{\rm MS}(f_{\rho})$ yields a two-term vector-field bound featuring a statistical rate in $\ell$ and a deterministic vLMM bias in $\mathbb{E}[H^{2p+2}]$, scaled by $c_{\rm obs}$.
V-C From Flow FS-HP Bound to Vector‑Field FS-HP Bound
With flow label $Y\in\mathcal{X}$ and $\rho((x,h),y)=\rho(y\mid x,h)\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}(x,h)$, the regression function is:
$\Phi^{h}_{\rho}(x)$ $\coloneqq\mathbb{E}[Y\mid X=x,H=h]$
$=\int_{\mathcal{X}}yd\rho(y\mid x,h).$
For a flow predictor $\Phi\in\mathcal{W}$,
$\mathcal{E}(\Phi)$ $\coloneqq\int_{\mathcal{Z}\times\mathcal{X}}\|\Phi^{h}(x)-y\|^{2}d\rho((x,h),y)$ (9)
$=\|\Phi-\Phi_{\rho}\|_{\mathcal{W}}^{2}+\mathcal{E}(\Phi_{\rho}).$
Thus minimizing $\mathcal{E}$ is equivalent to approximating $\Phi_{\rho}\in\mathcal{W}$ [4, Def. 1]. The choice of the estimated flow $\widehat{\Phi}\coloneqq\Phi_{z,y}^{\lambda_{\ell}}$ is inferred from data ${(z_{i},y_{i})}_{i=1}^{\ell}$ and induced from inductive bias encoded in $\mathcal{H}_{\mathcal{V}}$ through a kernel $k_{\mathcal{V}}$. Therefore, the empirical flow without vLMM modeling error is:
$\widehat{\Phi}\coloneqq\Phi_{z,y}^{\lambda_{\ell}}=\Phi(\ \cdot\ ;f_{z,y}^{\lambda_{\ell}}).$
Notice that the excess risk $\mathcal{E}(\widehat{\Phi})-\mathcal{E}(\Phi_{\rho})=\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}$ comes directly from (9), and similarly for $\mathcal{E}(\widehat{f})-\mathcal{E}(f_{\rho})$. Under the state design law $\rho_{\mathcal{X}}^{\mathrm{design}}$, flow and vector field excess risk are probabilistically bounded as follows.
Theorem 1 (Flow FS-HP bound).
Suppose $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ and $r\leq\mu$, which is the qualification $\mathsf{g}_{\lambda}$. Let $\beta=\max\{1,2\mu\}$. For $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\mathcal{E}(\widehat{\Phi})-\mathcal{E}(\Phi_{\rho})\leq\log\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}},$ (10)
for some finite positive constant $C_{\rm flow}$.
Proof.
[4, Thm. 2] applies unchanged to $\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}$. ∎
Theorem 2 (Vector-field FS-HP bound).
Suppose $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ and $r\leq\mu$, which is the qualification $\mathsf{g}_{\lambda}$. For $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\mathcal{E}(\widehat{f})-\mathcal{E}(f_{\rho})\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}}\log\frac{4}{\eta}\ell^{-\frac{2r}{2r+\beta}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}}\ \mathbb{E}[H^{2p+2}].$ (11)
for some finite positive constant $c_{\rm obs}$, $C_{\rm fit}$, and $C_{\rm bias}$.
Proof.
Let $\widehat{f}\coloneqq f_{z,y}^{\lambda_{\ell}}$ be the vector‑field estimator. Then:
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}$ $\leq\frac{1}{c_{\rm obs}^{2}}\ \|\mathsf{B}(\widehat{f}-f_{\rho})\|_{\mathcal{W}}^{2}$ (12)
$\leq\frac{1}{c_{\rm obs}^{2}}(\|\underbrace{\mathsf{B}\widehat{f}-\mathsf{Y}_{\alpha}\widehat{\Phi}}_{\Delta_{\rm MS}(\widehat{f})}\|_{\mathcal{W}}^{2}\ +\|\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})\|_{\mathcal{W}}^{2}$
$\qquad+\|\underbrace{\mathsf{Y}_{\alpha}\Phi_{\rho}-\mathsf{B}f_{\rho}}_{\Delta_{\rm MS}(f_{\rho})}\|_{\mathcal{W}}^{2})$ (13)
$\leq\frac{1}{c_{\rm obs}^{2}}\ \|\mathsf{J}+\mathsf{A}\|^{2}_{\mathcal{L}(\mathcal{W})}\big(\log\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}}\big)$
$\qquad+\frac{1}{c_{\rm obs}^{2}}\big(C^{2}_{{\rm LTE},\widehat{f}}+C^{2}_{{\rm LTE},f_{\rho}}\big)\mathbb{E}[H^{2p+2}],$ (14)
with $\mathcal{L}(\mathcal{W})$ the space of bounded linear operators $\mathcal{W}\to\mathcal{W}$.
First, inequality (12) follows immediately from the Lemma 1. Indeed, by definition $c_{\rm obs}^{2}=\lambda_{\min}(\mathsf{B}^{*}\mathsf{B})$, so for any $v\in\mathcal{V}$, $c_{\rm obs}^{2}\|v\|_{\mathcal{V}}^{2}\leq\|\mathsf{B}v\|_{\mathcal{W}}^{2}$; applying this with $v=\widehat{f}-f_{\rho}$ yields:
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}\leq\frac{1}{c_{\rm obs}^{2}}\|\mathsf{B}(\widehat{f}-f_{\rho})\|_{\mathcal{W}}^{2},$
which is exactly (12).
Next, to obtain (13), we add and subtract $\mathsf{Y}_{\alpha}\widehat{\Phi}$ and $\mathsf{Y}_{\alpha}\Phi_{\rho}$ inside the norm, and then apply the triangle inequality in $\mathcal{W}$ together with the linearity of the label map $\mathsf{Y}_{\alpha}=\mathsf{J}+\mathsf{A}$ as defined in (7). Using the identity:
$\mathsf{B}(\widehat{f}-f_{\rho})=\underbrace{\mathsf{B}\widehat{f}-\mathsf{Y}_{\alpha}\widehat{\Phi}}_{\Delta_{\rm MS}(\widehat{f})}+\underbrace{\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})}_{\text{flow mismatch}}+\underbrace{\mathsf{Y}_{\alpha}\Phi_{\rho}-\mathsf{B}f_{\rho}}_{\Delta_{\rm MS}(f_{\rho})},$
we directly arrive at the sum of the three squared $\mathcal{W}$-norms displayed in (13).
For the first term on the right-hand side of (14), i.e., the contribution of the flow mismatch, we use that $\mathsf{Y}_{\alpha}=\mathsf{J}+\mathsf{A}$ is a bounded linear operator on $\mathcal{W}$. In fact, $\mathsf{J}$ is an isometry and $\mathsf{A}$ is bounded by Assumption III.1. Hence, by sub-multiplicativity of the operator norm on $\mathcal{W}$,
$\|\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})\|_{\mathcal{W}}^{2}$ $=\|(\mathsf{J}+\mathsf{A})(\widehat{\Phi}-\Phi_{\rho})\|_{\mathcal{W}}^{2}$
$\leq\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}^{2}\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}.$
Invoking the flow FS–HP bound (Theorem 1) for $\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}$ then yields the first term appearing in (14).
For the second term on the right-hand side of (14), i.e., the contribution of the two vLMM residuals, we apply Lemma 2 to the first and third terms of (13). This gives $\|\Delta_{\rm MS}(\widehat{f})\|_{\mathcal{W}}^{2}\leq C_{{\rm LTE},\widehat{f}}^{2}\mathbb{E}[H^{2p+2}],$ and $\|\Delta_{\rm MS}(f_{\rho})\|_{\mathcal{W}}^{2}\leq C_{{\rm LTE},f_{\rho}}^{2}\mathbb{E}[H^{2p+2}],$ which together produce the second term in (14).
Finally, note that $\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}$ is bounded under Assumption III.1. Therefore we can absorb constants by setting $C_{\rm fit}\geq C_{\rm flow}\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}^{2}$ and $C_{\rm bias}\coloneqq C_{{\rm LTE},\widehat{f}}^{2}+C_{{\rm LTE},f_{\rho}}^{2}$. Substituting these into (14) completes the proof of (11). ∎
An immediate insight from (11) is that the excess risk compose of two parts: (i) the statistic fitting error term $\log(4/\eta)\ell^{-{2r}/{(2r+\beta)}}$ that scales down as number of sampling data increase; (ii) the discretization error term $\mathbb{E}[H^{2p+2}]$ that scales down as the average step size goes to zero. The discretization error is independent of $\ell$, therefore, Theorem 2 predicts that we need to control both $\ell$ and $H$ to reduce the excess risk of $\widehat{f}$. However, there are some caveats to be addressed regarding the $\mathsf{B}$-observability constant $c_{\rm obs}$.
VI Conclusion
We introduced Spectral Flow Learning (SFL), a framework that learns dynamics in a windowed flow space and then lifts guarantees to the vector field. The key modeling device is a lag-linear label operator that aggregates lagged Koopman actions through variable-step linear multistep (vLMM) windows. In this space, we deploy spectral regularization with qualification-controlled filters, yielding finite-sample, high-probability bounds for the flow predictor. A multistep observability inequality induced by the Koopman-lag forcing operator links flow error to vector-field error, and a uniform local-truncation bound controls the deterministic discretization bias due to vLMM. Together these ingredients produce a two-term vector-field FS-HP bound that separates a statistical rate, $\mathcal{O}\left(\ell^{-\frac{2r}{2r+\beta}}\right)$, from an explicit vLMM bias governed by $\mathbb{E}[H^{2p+2}]$, revealing a principled step-sample tradeoff. The analysis applies across AB/AM/BDF families under bounded step ratios and order-$p$ consistency, and it is compatible with physics-informed vvRKHS designs.